OpenStudy (anonymous):

One more additional question How many terms of the series 9 + 0.1 + 0.01 + 0.001 + ... must be taken so that the sum of infinity exceeds the\[n^{th}\] partial sum by \[\frac{ 1 }{ 9\times10^{7} }\]

OpenStudy (anonymous):

choices A) 6 B) 7 C)9 D)11 E) None

OpenStudy (anonymous):

I think its 1+0.1+0.01+0.001+...

OpenStudy (anonymous):

nope, It just like it.

OpenStudy (anonymous):

What is partial sum

OpenStudy (anonymous):

the sum of the series up to n terms

OpenStudy (anonymous):

Oh..... I just got the question/

OpenStudy (anonymous):

7

OpenStudy (anonymous):

U have to take less than 7.95 terms

OpenStudy (anonymous):

You are triked

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Isnt Sum of infinity =82/9

OpenStudy (anonymous):

Read the question carefully

OpenStudy (anonymous):

Oh forgot to add 1

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yape

OpenStudy (anonymous):

\[S _{\infty} = S_{n} + \frac{ 1 }{ 9\times10^{7} }\] \[9+\frac{ \frac{ 1 }{ 10 }(1-\frac{ 1 }{ 10^{\infty} }) }{ 1-\frac{ 1 }{ 10 } } =9+\frac{ \frac{ 1 }{ 10 }(1-\frac{ 1 }{ 10^{n-1} }) }{ 1-\frac{ 1 }{ 10 } }+\frac{ 1 }{ 9\times10^{7} }\]

OpenStudy (anonymous):

So, Is the answer none.

OpenStudy (anonymous):

yes